# A Foundation Model for General Moving Object Segmentation in Medical Images

### v1

@@ 5.2 Conclusion
In this work, we propose iMOS, a foundation model for general moving object segmentation in medical images. iMOS can automatically segment the targets of the remaining frames of the video based on the segmentation mask of the first frame provided by the video. At the same time, we regard each slice of medical 3D volume as each frame of the video, so iMOS can also perform segmentation on medical 3D volume. We have demonstrated its validity and stability on medical images through a series of experiments. iMOS provides technical accumulation for auxiliary medical video or 3D volume annotation tasks and surgical navigation.

### v5

@@ 5 Conclusion
In this work, we propose iMOS, a foundation model for general moving object segmentation in medical images. iMOS can automatically segment the remaining targets based on the manually-annotated mask of the first frame/slice in video/volume. We have demonstrated its validity and generalizability through extensive experiments. We believe that iMOS can provide technical accumulation for auxiliary medical video or 3D volume annotation tasks, and boost the development of the medical image analysis community.
6 Compliance with ethical standards
The research was conducted using medical images in open access. Ethical approval was not required as confirmed by the license attached with the open access data.
7 Acknowledgements
This work was supported by the grant from National Natural Science Foundation of China (Nos. 62101343, 62171290), Shenzhen-Hong Kong Joint Research Program (No. SGDX20
201103095613036), Science and Technology Planning Project of Guangdong Province, China (2023A0505020002).